# Communication Efficient LLM Pre-training with SparseLoCo

### v1

@@ Abstract
Communication-efficient distributed training algorithms have received considerable interest recently due to their benefits for training Large Language Models (LLMs) in bandwidth-constrained settings, such as across data centers and over the internet. Despite reducing communication frequency, these methods still typically require communicating a full copy of the model’s gradients—resulting in a communication bottleneck even for cross-datacenter links. Furthermore, they can slightly degrade performance compared to a naive AdamW DDP baseline. While quantization and error feedback are often applied to reduce the pseudo-gradient’s size, in the context of LLM pre-training, existing approaches have been unable to additionally leverage sparsification and have obtained limited quantization. In this work, we introduce SparseLoCo, a communication-efficient training algorithm for LLMs that effectively leverages sparsification and quantization to reach extreme compression ratios of up to 1–3% sparsity and 2-bit quantization while outperforming full-precision DiLoCo. Our key observations are that outer momentum can be locally approximated by an error feedback combined with aggressive sparsity and that sparse aggregation can actually improve model performance. We empirically demonstrate in a range of communication-constrained LLM training settings that SparseLoCo provides significant benefits in both performance and communication cost.
@@ 1 Introduction
Frontier language models pre-trained on internet-scale data have led to considerable breakthroughs in recent years. However, due to their growing parameter counts, effectively training these models across expensive datacenter hardware while retaining efficiency—a central goal due to the resources spent on these runs—is becoming increasingly challenging. On the other hand, due to the increasing availability of globally distributed computational infrastructure across the world, the pre-training of large-scale models over the internet has recently garnered special attention (Jaghouar et al., 2024). Similar to training over the internet, pre-training across multiple data centers requires mitigating the large communication overhead incurred by aggregating updates between workers.
In the context of LLM pre-training, several approaches have been proposed to reduce data-parallel communication cost. Among them are DiLoCo (Douillard et al., 2023b), a variant of LocalSGD (Stich, 2018; Reddi et al., 2020), as well as methods compressing communicated tensors and leveraging error feedback (Peng et al., 2024; Ahn & Xu, 2025; Wang et al., 2023) to mitigate information loss. These techniques have complementary advantages of (1) reducing the communication frequency and (2) reducing the size of communicated messages. Combining the two is potentially advantageous for bandwidth-constrained settings like training over the internet or across data centers. However, existing works focused on LLM pre-training (Charles et al., 2025) do not take full advantage of compression schemes.
Indeed, combining these approaches raises the challenge of how to incorporate error feedback with an outer momentum, which is known to be important for DiLoCo’s performance. Our key observation is that when aggressive sparsification is combined with error feedback on DiLoCo pseudo-gradients, two effects emerge: (a) error feedback naturally acts as a local approximation of outer momentum, and (b) sparse aggregation is induced on the pseudo-gradients, a property recently shown in model merging contexts to improve performance (Yadav et al., 2023; Davari & Belilovsky, 2024). Building on this, we introduce SparseLoCo, which replaces global outer momentum with a single error feedback accumulator, thereby unifying infrequent and sparsified communication. This enables aggressive sparsification and quantization of pseudo-gradients, while outperforming full-communication DiLoCo and frequency-compressed baselines.
@@
We demonstrate empirically that outer momentum in DiLoCo can be replaced with a local approximation, which we link to with error feedback.
@@
Through our experiments, we demonstrate that SparseLoCo can significantly reduce the communication volume compared to existing LLM training methods (e.g., DiLoCo and DeMo), while simultaneously outperforming them.
@@ 2 Related Work
In the federated learning literature, communication efficiency has been a central focus from the outset, as participating clients often operate over highly constrained and heterogeneous networks. A canonical example is Federated Averaging (FedAvg) (Konečnỳ et al., 2016; McMahan et al., 2017), which reduces communication frequency by performing multiple local updates before averaging model parameters. Other works explore compressed updates through sketching or quantization in the context of federated learning (Rothchild et al., 2020; Reisizadeh et al., 2020). Beyond reducing communication overhead, numerous approaches such as SCAFFOLD (Karimireddy et al., 2020) and FedProx (Li et al., 2020) address the unique challenge of data heterogeneity—where each client’s dataset may follow a different distribution—by introducing control variates or proximal terms to stabilize convergence. While our work shares federated learning’s emphasis on reducing communication overhead, it differs fundamentally in scope: we focus on large-scale pre-training of LLMs in settings with homogeneous data partitions (e.g., sharded web-scale corpora), where heterogeneity-mitigation strategies are unnecessary.
LocalSGD and extensions to LLM training
@@
Local Stochastic Gradient Descent (LocalSGD) (Stich, 2018) is a widely studied approach for reducing communication in distributed training by allowing workers to perform multiple local updates before synchronizing. Stich (2018) formally introduced the method and proved its convergence, while Lin et al. (2018) highlighted that LocalSGD can lead to improved generalization compared to simply increasing the batch size. Extensions of LocalSGD include SlowMo (Wang et al., 2019), which incorporates a slow outer momentum to stabilize training in datacenter-style environments—while still using SGD as the inner optimizer—and meta-learning approaches (Joseph et al., 2025) that adapt the aggregation function for improved performance. However, these approaches were not shown to scale well to pre-training in Ortiz et al. (2021). More recently, DiLoCo (Douillard et al., 2023b) adapted the LocalSGD framework to Large Language Model (LLM) pre-training, demonstrating that replacing the inner optimizer with AdamW can yield substantial benefits. Our work builds upon this line of research by enabling aggressive sparsification of the communicated pseudo-gradients in a LocalSGD-style framework, something that prior methods have not achieved while maintaining or improving upon state-of-the-art LLM training performance. Finally, Douillard et al. (2025) and Fournier et al. (2024) consider communicating a small subset of model parameters more frequently instead of communicating a message the size of the model infrequently by allowing the models to desynchronize while still remaining relatively close to each other (in terms of, e.g., consensus distance). SparseLoCo, on the other hand, maintains the benefit of infrequent communication while communicating a small-sized message and maintaining model synchronization.
Error Feedback and Compressed Updates
@@
Error feedback (EF) has been extensively studied, particularly from a theoretical perspective, as a means to compensate for the information loss introduced by various gradient compression methods (Seide et al., 2014; Karimireddy et al., 2019; Stich & Karimireddy, 2019). It has been combined with various compression techniques, including quantization, sparsification (Shi et al., 2019), and low-rank approximation in (Vogels et al., 2019; Ahn & Xu, 2025), and applied to LLMs in recent works (Wang et al., 2023; Peng et al., 2024). EF21-SGDM (Fatkhullin et al., 2023) analyzed how to combine error feedback with momentum, introducing a momentum-compatible variant that requires two accumulators and is largely focused on theoretical aspects and does not address the multi-iteration setting or the practical challenges of LLM pre-training. QSparseLocalSGD (Basu et al., 2019) is, to our knowledge, one of the few works that combines multi-iteration methods such as LocalSGD with error feedback, but its focus was on theoretical analysis with non-adaptive optimizers and without outer momentum. In contrast, our work targets the LLM pre-training regime and develops a method to combine aggressive compression and error feedback with an efficient approximation of outer momentum. DeMo (Peng et al., 2024) considers EF with DCT encoding and compression in the LLM setting, demonstrating it can achieve competitive performance, but without incorporating local updates or the ability to leverage adaptive optimizers. Similarly, CocktailSGD (Wang et al., 2023) uses error feedback with multiple compression operators in an LLM fine-tuning setting, yet does not explore the integration of local iteration methods. To the best of our knowledge, this is the first work to study the combination of these approaches in the context of LLMs and more generally in the context of modern variations of multi-iteration methods that have been shown to scale to pre-training.
@@ 3.1 Background and Notation
Consider the DiLoCo/FedOpt (Douillard et al., 2023a; Reddi et al., 2020) framework, which utilizes the following basic rule on each worker or replica to produce a pseudo-gradient at each outer step, , as follows:
Here, corresponds to the number of inner steps of the optimizer (typically AdamW), and is the number of replicas. DiLoCo, which corresponds to an instantiation of FedOpt with AdamW as the inner optimizer and outer (server) momentum using Nesterov (Dozat, 2016), is given as follows:
@@ 3.2 Local Outer Momentum
We first propose a variant of DiLoCo that utilizes a per-replica local outer momentum instead of the unified global momentum. The goal of this algorithm is to provide insight into how well the outer momentum can be locally approximated. We denote this variant of DiLoCo as DiLoCo-LOM (Local Outer Momentum):
Here, the outer momentum is updated locally, solely based on the local pseudo-gradient, while the final update is based on the average of the local momentum accumulators. Note that typical implementations of DiLoCo store the outer momentum locally on each replica, meaning that DiLoCo-LOM does not add any memory overhead compared to the global momentum variant.
Building up to SparseLoCo, we consider an additional method, denoted DiLoCo-LOM-Sub-, where the local momentums have their largest components removed at the end of each outer step:
Here, the local momentum has its largest-magnitude components removed. This allows us to study the impact of subtraction, used in error feedback, without sparsifying the pseudo-gradient.
@@ 3.3 SparseLoCo: Sparse Aggregation meets Local Outer Momentum
We now introduce SparseLoCo, which blends sparsification and error feedback in place of the local outer momentum. We consider error feedback, , applied to the pseudo-gradients, which we denote OuterEF:
Here, is the quantization function that also allows us to leverage the error feedback to further reduce the message size transmitted. When is sufficiently small, OuterEF closely approximates the local outer momentum in LOM, since only a few components will be subtracted from . On the other hand, unlike LOM and LOM-Sub-, SparseLoCo only aggregates quantized sparse vectors, drastically reducing the message size needed for communication. The full algorithm for SparseLoCo is given in Algorithm 1.
SparseLoCo uses a variant of the operation inspired by Peng et al. (2024) that divides tensors into discrete chunks and applies the within each chunk. This has three benefits compared to applying it at the full-tensor or global level: (a) the cost of naively storing indices for transmission is significantly reduced as each chunk’s index space is bounded. (b) applied to entire models or individual tensors can overemphasize correlated variables; thus, chunking can have benefits on performance as further discussed in Appendix B. (c) Finally, chunking can allow for more easily integrating tensor parallelism and FSDP, which often require sharding across tensors, thereby creating inefficiencies for operations over entire tensors or models.
@@ 4 Experiments
Our experiments use a 512M-parameter, LLaMA-style decoder-only transformer on DCLM (Li et al., 2024) using the LLaMA-2 tokenizer (Touvron et al., 2023). Following Hoffmann et al. (2022), we allocate a token budget equal to the model size. Our experimental protocol follows (Charles et al., 2025). Unless otherwise stated, we use a per-replica batch size with sequence length and replicas, yielding a global batch of M tokens per step. In SparseLoCo, we employ 2-bit quantization with a chunk size of . We report the hyperparameter sweep ranges, architectural details, and the best settings in Appendix D. As baselines, we include DiLoCo and DeMo—two strong communication-efficient methods (one using local iterations and the other using error feedback) for LLMs—as well as an AdamW baseline.
@@ 4.1 Building intuition with Local Outer Momentum
We first use the DiLoCo-LOM and DiLoCo-LOM-Sub- algorithms to empirically link the standard outer momentum in DiLoCo to the error feedback mechanism in SparseLoCo. In DiLoCo-LOM, each replica maintains a local outer momentum accumulator that is averaged only at synchronization. In DiLoCo-LOM-Sub-, we subtract the largest entries of the local momentum after each synchronization, isolating the effect of subtraction while keeping the communicated pseudo-gradients dense. As shown in Table 1, DiLoCo-LOM matches DiLoCo’s loss, and pruning of the largest accumulator entries has a negligible impact, whereas removing outer momentum entirely degrades performance.
To quantify how closely local outer momentum tracks the target global outer momentum in DiLoCo, we maintain a reference global accumulator and, over the first outer steps, compute the cosine similarity between this reference and each replica’s local accumulator at corresponding steps. The average similarity is for DiLoCo-LOM and for DiLoCo-LOM-Sub- (25%), indicating that sparsified local accumulators remain a strong directional proxy for the global momentum and supporting our interpretation of SparseLoCo’s error feedback state as a local approximation of DiLoCo’s outer momentum.
| Method | Loss |
| DiLoCo | 2.760 |
| DiLoCo w.o. outer momentum | 2.868 |
| DiLoCo-LOM | 2.759 |
| DiLoCo-LOM-Sub- - 25% | 2.761 |

### v2

@@ Abstract
Communication-efficient distributed training algorithms have received considerable interest recently due to their benefits for training Large Language Models (LLMs) in bandwidth-constrained settings, such as across datacenters and over the internet. Despite reducing communication frequency, these methods still typically require communicating a full copy of the model’s gradients—resulting in a communication bottleneck even for cross-datacenter links. Furthermore, they can slightly degrade performance compared to a naive AdamW DDP baseline. While quantization is often applied to reduce the pseudo-gradient’s size, in the context of LLM pre-training, existing approaches have been unable to additionally leverage sparsification and have obtained limited quantization. In this work, we introduce SparseLoCo, a communication-efficient training algorithm for LLMs that effectively leverages error feedback with sparsification and 2-bit quantization to reach extreme sparsity as low as 1–3% while outperforming full-precision DiLoCo. Our key observations are that outer momentum can be locally approximated by an error feedback accumulator combined with aggressive sparsity, and that sparse aggregation can actually improve model performance. We empirically demonstrate in a range of communication-constrained LLM training settings that SparseLoCo provides significant benefits in both performance and communication cost.
@@ 1 Introduction
Frontier language models pre-trained on internet-scale data have led to considerable breakthroughs in recent years. However, due to their growing parameter counts, effectively training these models across expensive datacenter hardware while retaining efficiency—a central goal due to the resources spent on these runs—is becoming increasingly challenging. On the other hand, due to the increasing availability of globally distributed computational infrastructure across the world, the pre-training of large-scale models over the internet has recently garnered increasing interest (intellect1). Similar to training over the internet, pre-training across multiple datacenters requires mitigating the large communication overhead incurred by aggregating updates between workers.
In the context of LLM pre-training, several approaches have been proposed to reduce data-parallel communication cost. Among them are DiLoCo (DiLoCo), a variant of LocalSGD (stich2018local; reddi2020adaptive), as well as methods compressing communicated tensors and leveraging error feedback (peng2024decoupled; ahn2025dion; pmlr-v202-wang23t) to mitigate information loss. These techniques have complementary advantages of (1) reducing the communication frequency and (2) reducing the size of communicated messages. Combining the two is potentially advantageous for bandwidth-constrained settings like training over the internet or across datacenters. However, existing works focused on LLM pre-training using DiLoCo (scalingdiloco) do not take full advantage of compression schemes.
Indeed, combining these approaches raises the challenge of how to incorporate error feedback with an outer momentum, which is known to be important for DiLoCo’s performance. Our key observation is that when aggressive sparsification is combined with error feedback on DiLoCo pseudo-gradients, two effects emerge: (a) error feedback naturally acts as a local approximation of outer momentum, and (b) sparse aggregation is induced on the pseudo-gradients, a property recently shown in model merging contexts to improve performance (yadav2023tiesmerging; davari2024model). Building on this, we introduce SparseLoCo, which replaces global outer momentum with a single error feedback accumulator, thereby unifying infrequent and sparsified communication. This enables aggressive sparsification and quantization of pseudo-gradients, while outperforming full-communication DiLoCo and frequency-compressed baselines.
@@
We demonstrate that DiLoCo’s outer momentum can be replaced with a local momentum, which we link to with error feedback on pseudo-gradients.
@@
Through our extensive experiments, we demonstrate that SparseLoCo can significantly reduce the communication volume compared to existing LLM training methods (e.g., DiLoCo and DeMo), while simultaneously outperforming them.
@@ 2 Related Work
Federated Learning In the federated learning literature, communication efficiency has been a central focus from the outset, as participating clients often operate over highly constrained and heterogeneous networks. A canonical example is Federated Averaging (FedAvg) (konevcny2016federated; mcmahan2017communication), which reduces communication frequency by performing multiple local updates before averaging model parameters. Other works explore compressed updates through sketching or quantization in the context of federated learning (fetchsgd; qfed). Beyond reducing communication overhead, numerous approaches such as SCAFFOLD (karimireddy2020scaffold) and FedProx (li2020federated) address the unique challenge of data heterogeneity—where each client’s dataset may follow a different distribution—by introducing control variates or proximal terms to stabilize convergence. Related to our work mitchell2022optimizing consider pseudo-gradient compression in the FL setting. While our work shares federated learning’s emphasis on reducing communication overhead, it differs fundamentally in scope: we focus on large-scale pre-training of LLMs in settings with homogeneous data partitions (e.g., sharded web-scale corpora), where heterogeneity-mitigation strategies are unnecessary while achieving performance that can match standard data parallel schemes at equivalent FLOPs is the paramount douillard2023diloco.
@@
LocalSGD and extensions to LLM training Local Stochastic Gradient Descent (LocalSGD) (stich2018local) is a widely studied approach for reducing communication in distributed training by allowing workers to perform multiple local updates before synchronizing. stich2018local formally introduced the method and proved its convergence, while lin2018local highlighted that LocalSGD can lead to improved generalization compared to simply increasing the batch size. Extensions of LocalSGD include SlowMo (wang2019slowmo), which incorporates a slow outer momentum to stabilize training in datacenter-style environments—while still using SGD as the inner optimizer—and meta-learning approaches (metalearningcomm) that adapt the aggregation function for improved performance. However, these approaches were not shown to scale well to pre-training in ortiz2021trade. More recently, DiLoCo (DiLoCo) adapted the LocalSGD framework to Large Language Model (LLM) pre-training, demonstrating that replacing the inner optimizer with AdamW and a nesterov momentum outer optimizer can yield substantial benefits. Our work builds upon this line of research by enabling aggressive sparsification of the communicated pseudo-gradients in a LocalSGD-style framework, something that prior methods have not achieved while maintaining or improving upon state-of-the-art LLM training performance. Finally, streamingdiloco and fournier2024wash consider communicating a small subset of model parameters more frequently instead of communicating a message the size of the model infrequently by allowing the models to desynchronize while still remaining relatively close to each other (in terms of, e.g., consensus distance). SparseLoCo, on the other hand, maintains the benefit of infrequent communication while communicating a small-sized message and maintaining model synchronization.
@@
Error Feedback and Compressed Updates Error feedback (EF) has been extensively studied, particularly from a theoretical perspective, as a means to compensate for the information loss introduced by various gradient compression methods (seide20141bit; karimireddy2019error; stich2019error). It has been combined with various compression techniques, including quantization, sparsification (shi2019topk), and low-rank approximation in (powersgd; ahn2025dion). In the single local step setting it has been applied to LLMs in recent works (pmlr-v202-wang23t; peng2024decoupled; zhaoseparate). EF21-SGDM (ef21) analyzed how to combine error feedback with momentum, introducing a momentum-compatible variant that requires two accumulators and is largely focused on theoretical aspects and does not address the multi-iteration setting or the practical challenges of LLM pre-training. QSparseLocalSGD (basu2019qsparselocalsgd) is, to our knowledge, one of the few works that combines multi-iteration methods such as LocalSGD with error feedback, but its focus was on theoretical analysis with non-adaptive optimizers and without outer momentum which is crucial to high performance in the LLM setting. In contrast, our work targets the LLM pre-training regime and develops a method to combine aggressive compression and error feedback with an efficient approximation of outer momentum. DeMo (peng2024decoupled) considers EF with DCT encoding and compression in the LLM setting, demonstrating it can achieve competitive performance, but without incorporating local updates or the ability to leverage adaptive optimizers. Similarly, CocktailSGD (pmlr-v202-wang23t) uses error feedback with multiple compression operators in an LLM fine-tuning setting, yet does not explore the integration of local iteration methods. Our work studies the combination of these approaches in the context of LLMs and more generally in the context of modern variations of multi-iteration methods that have been shown to scale to pre-training.
@@ 3.1 Background and Notation
Consider the DiLoCo/FedOpt (douillard2023diloco; reddi2020adaptive) framework, which utilizes the following basic rule on each worker or replica to produce a pseudo-gradient at each outer step, , as follows:
Here, corresponds to the number of inner steps of the optimizer (typically AdamW), and is the number of replicas. DiLoCo, which corresponds to an instantiation of FedOpt with AdamW as the inner optimizer and outer (server) momentum using Nesterov (dozat2016incorporating), is given as follows:
@@ 3.2 Local Outer Momentum
We first propose a variant of DiLoCo that utilizes a per-replica local outer momentum instead of the unified global momentum. The goal of this algorithm is to provide insight into how well the outer momentum can be locally approximated. We denote this algorithm DiLoCo-LOM (Local Outer Momentum):
Here, the outer momentum is updated locally, solely based on the local pseudo-gradient, while the final update is based on the average of the local momentum accumulators . Note that typical implementations of DiLoCo store the outer momentum locally on each replica, meaning that DiLoCo-LOM does not add any memory overhead compared to the global momentum variant. We show that the DiLoCo-LOM update exactly matches the DiLoCo update in Appendix I.
Building up to SparseLoCo, we consider an additional method, denoted DiLoCo-LOM-Sub-, where the local momenta have their largest components removed at the end of each outer step:
This allows us to study the impact of subtraction, used in error feedback, without sparsifying the pseudo-gradient.
@@ 3.3 SparseLoCo: Sparse Aggregation meets Local Outer Momentum
We now introduce SparseLoCo, which blends sparsification and error feedback in place of the local outer momentum. We consider error feedback, , applied to the pseudo-gradients, which we denote as OuterEF:
Here, is the quantization function which allows further compression of the selected values. When is sufficiently small, OuterEF closely approximates the local outer momentum in LOM, since only a few components will be subtracted from . On the other hand, unlike LOM and LOM-Sub-, SparseLoCo only aggregates quantized sparse vectors, drastically reducing the message size needed for communication. The full algorithm for SparseLoCo is given in Algorithm 1.
SparseLoCo uses a variant of the operation inspired by 9546514; peng2024decoupled that divides tensors into discrete chunks and applies the within each chunk. This has three benefits compared to applying it at the full-tensor or global level: (a) the cost of naively storing indices for transmission is significantly reduced as each chunk’s index space is bounded. (b) applied to entire models or individual tensors can overemphasize correlated variables; thus, chunking can have benefits on performance as further discussed in Appendix B. (c) Finally, chunking can allow for more easily integrating tensor parallelism and FSDP, which often require sharding across tensors, thereby creating inefficiencies for operations over entire tensors or models.
@@ 4 Experiments
Our experiments use 178M-, 512M-, and 2B-parameter LLaMA-style decoder-only transformer on DCLM (dclm) using the LLaMA-2 tokenizer (touvron2023llama). Following hoffmann2022training, we allocate a token budget equal to the model size. Our experimental protocol follows (scalingdiloco). Unless otherwise stated, our main results are reported on the 512M-parameter model with workers, per-worker batch size , and sequence length , yielding a global batch of M tokens per step. For SparseLoCo, we employ 2-bit quantization with a chunk size of (Non-overlapping square grids for 2D parameters). We apply a short error feedback (OuterEF) freeze, where the error feedback is not utilized for the first of the outer steps to improve training stability and performance (ablated in Table 11 in Appendix E). We further study scaling across model sizes (178M; Appendix F, 2B; Table 5), and number of workers in Table 4, and Appendix Tables 12, and 13. We report the hyperparameter sweep ranges, architectural details, and selected configurations in Appendix H. As baselines, we include DiLoCo and DeMo—two strong communication-efficient methods (one using local iterations and the other using error feedback) for LLMs—as well as a DDP AdamW baseline.
@@ 4.1 Building intuition with Local Outer Momentum
Method Loss DiLoCo 2.760 DiLoCo w.o. outer momentum 2.868 DiLoCo-LOM 2.759 DiLoCo-LOM-Sub- - 25% 2.761
We first use the DiLoCo-LOM and DiLoCo-LOM-Sub- algorithms to empirically link the standard outer momentum in DiLoCo to the error feedback mechanism in SparseLoCo. In DiLoCo-LOM, each replica maintains a local outer momentum accumulator that is averaged only at synchronization. In DiLoCo-LOM-Sub-, we subtract the largest entries of the local momentum after each synchronization, isolating the effect of subtraction while keeping the communicated pseudo-gradients dense. As shown in Table 1, DiLoCo-LOM matches DiLoCo’s loss—consistent with Proposition 1 (Appendix I)—and pruning of the largest accumulator entries has a negligible impact, whereas removing outer momentum entirely degrades performance.
To quantify how closely local outer momentum tracks the target global outer momentum in DiLoCo, we maintain a reference global accumulator and, over the first outer steps, compute the cosine similarity between this reference and each replica’s local accumulator at corresponding steps. The average similarity between individual local accumulators and the global reference accumulator is for DiLoCo-LOM-Sub- (25%), indicating that removing the top components from local accumulators each outer step still allows them to remain a strong directional proxy for the global momentum and supporting our interpretation of SparseLoCo’s error feedback state as a local approximation of DiLoCo’s outer momentum.
| Method | Density | Loss | Pseudo-Grad Size | # of Syncs | Quantization |
| AdamW DDP | 100% | 2.69 | 1.02 GB | 2445 | 16-bit |
| DiLoCo (H=15) | 100% | 2.76 | 512.40 MB | 163 | 8-bit |
| DeMo | 0.78% | 2.83 | 10.01 MB | 2445 | 8-bit |
| DeMo | 3.12% | 2.86 | 40.03 MB | 2445 | 8-bit |
| SparseLoCo (H=15) | 0.78% | 2.79 | 4.25 MB | 163 | 2-bit |
| SparseLoCo (H=15) | 3.12% | 2.70 | 17.01 MB | 163 | 2-bit |